# Magnetic disk winds in protoplanetary disks:

### v2

@@ 2.1 Disk evolution with FEOSAD
In this section, we summarize the underlying MHD model that is used for the evolution of the PPD in the thin-disk limit. The magnetic disk winds described in Section 2.2 are added to this base model in a modular fashion. The hydrodynamic simulations are conducted using the code Formation and Evolution Of a Star And its circumstellar Disk (FEOSAD, Vorobyov et al., 2018a, 2020b). The numerical implementation of hydrodynamic algorithms is based on the operator-split procedure similar in methodology to the ZEUS code (Stone & Norman, 1992). We take into account the self gravity of the disk, co-evolution of the coupled two-part dust component, calculation of magnetic fields in the flux-freezing approximation, and a dead zone modeled with an adaptive-$\alpha$ formulation, where the effective $\alpha$ is derived from ionization-recombination balance equations. Here we describe the minimum set of equations, while a more detailed description of the model can be obtained in Vorobyov et al. (2020b) and references therein.
The equations of mass continuity, conservation of momentum, and energy transport for the gas component in the thin disk limit can be described as follows.
@@
The energy transport equation (Eq. (3)) pertains to the evolution of the internal energy per surface area, $e$. The disk midplane temperature ($T_{\rm mp}$), which is assumed equal for the dust and gas components, is obtained by solving this energy balance equation. The heating ($\Gamma$) and cooling ($\Lambda$) rates are based on the analytical solution of the radiation transfer equations in the vertical direction (Dong et al., 2016). The heating function takes into account the irradiation of the disk surface from the stellar as well as background irradiation and radiation cooling by dust grains is also considered (Vorobyov et al., 2018a). The irradiation from the central star is calculated using incidence angle of radiation arriving at the disk surface, which in turn is obtained using the local scale height and a fixed flaring angle (Vorobyov & Basu, 2010b; Dong et al., 2016). These calculations neglect the self-shielding of the disk, which may arise in the innermost regions due to formation of concentric rings. However, this effect is considered to be insignificant, as viscous heating dominates the energy balance in the midplane of this inner disk region. The resulting model has a flared structure, wherein the disk vertical scale height increases with radial distance. Both the disk and the infalling envelope receive a fraction of the irradiation energy from the central protostar, both in terms of accretion and photospheric luminosity (Vorobyov & Basu, 2009). The optical depths in the calculations are proportional to the total dust surface density, however, they do not take into account the dust growth. The Planck and Rosseland mean opacities needed for these calculations are taken from Semenov et al. (2003). The term involving viscous stress tensor accounts for the viscous heating, which typically dominates the inner disk and becomes significant for an extended region during outbursts (Vorobyov & Basu, 2010b; Kadam et al., 2020). The last term in Eq. (3) corresponds to the internal energy of matter carried away by the magnetic winds.
@@
where $\Sigma_{\rm d,sm}$ and $\Sigma_{\rm d,gr}$ are the surface densities of small and grown dust, respectively. The term $u$ denotes the planar components of the grown dust velocity and $\dot{\Sigma}_{\rm w,d}$ accounts for the small dust blown away with the disk winds. Since the winds originate in the upper layers of the disk surface, we assume that they do not carry grown dust. $S$ is the rate of conversion from small to grown dust per unit surface area, which is a function of the maximum size of the dust $(a_{\rm max})$. The dust is assumed to mix vertically with the gas, which is a reasonable approximation for a young disk evolving under gravitational and viscous torques. $S(a_{\rm max})$ is derived from the assumption that the size distributions of both the dust populations can be described by a power law with an exponent of $-3.5$. The discontinuity in the dust size distribution at $a_{*}$ is assumed to get smoothed out, which implies the dominant role of dust growth as compared to the dust flow (Molyarova et al., 2021). The evolution of the maximum size of the grown dust can be expressed as an advection equation modified for the presence of a source term
@@
where $v_{\rm frag}$ is the fragmentation velocity, $c_{\rm s}$ is the sound speed and $\alpha_{\rm turb}$ corresponds to the turbulence parameter (Birnstiel et al., 2012), wherein the possible contribution from gravitoturbulence is not considered (see, e.g., Vorobyov et al., 2023). We set the fragmentation velocity to a canonical value of 3 ${\rm m\,s^{-1}}$, although there is significant ambiguity due to sticking properties of different grain compositions, e.g., bare silicate versus icy grains (Blum & Wurm, 2008; Wada et al., 2009; Steinpilz et al., 2019). When $a_{\rm max}$ exceeds $a_{\rm frag}$, the growth rate $\cal{D}$ is set to zero and $a_{\rm max}$ is set equal to $a_{\rm frag}$, which guarantees that the particle size is capped at $a_{\rm frag}$. We note that in our approach, the dust growth is limited to keep the size of dust particles within the Epstein regime, so that the back-reaction term is expressed as
@@
The viscosity in a protoplanetary disk is thought to originate primarily due to the turbulence generated by MRI (Balbus & Hawley, 1991; Turner et al., 2014). As the cosmic rays externally penetrate the disk surface, the inner disk ($\lesssim 10$ au) accretes through the MRI-active surface layers, while a magnetically dead zone is formed at the disk midplane (Gammie, 1996; Simon et al., 2013). Although there may be quantitative uncertainty, this picture of layered accretion is considered generally valid, even in the presence of disk winds (Armitage & Kley, 2019). We model the magnetically dead zone using an effective and adaptive Shakura & Sunyaev (1973) $\alpha$ parameter (Bae et al., 2014; Kadam et al., 2019)-
@@
where $\Sigma_{{\rm MRI}}$ is the gas surface density of the MRI-active layer and $\Sigma_{{\rm dz}}$ is that of the magnetically dead layer, so that $\Sigma_{\rm g}=\Sigma_{{\rm MRI}}+\Sigma_{{\rm dz}}$. Here, $\alpha_{\rm MRI}$ and $\alpha_{{\rm dz}}$ correspond to the strength of the turbulence in the MRI-active layer and the dead zone, respectively. The $\Sigma_{\rm MRI}$ is obtained from the ionization fraction ($x$), which in turn is determined by solving the ionization balance equation
@@
where $\xi$ is the ionization rate, $\alpha_{\rm{r}}$ is radiative recombination rate, $n_{\rm n}$ is the number density of neutrals, and $\alpha_{\rm{d}}$ is the total rate of recombination on to the dust grains (Dudorov & Khaibrakhmanov, 2014; Dudorov & Sazonov, 1987; Balduin et al., 2023). We consider the ionization by cosmic rays and radionuclides, however, effects of stellar far-ultraviolet and X-ray radiation are neglected, since their contribution is typically significantly smaller in the inner disk and limited to the upper disk atmosphere, not resolved in our thin-disk model (Igea & Glassgold, 1999; Bergin et al., 2007; Perez-Becker & Chiang, 2011).
In the simulations, $\alpha_{\rm MRI}$ is set to the canonical value of $0.01$, while lower values of $10^{-3}$ and $10^{-4}$ are also explored, the latter being consistent with recent observations of low viscosity in PPDs (Flaherty et al., 2015; Dullemond et al., 2018; Muro-Arena et al., 2018; Flaherty et al., 2020; Franceschi et al., 2023). 3D MHD simulations of FU Orionis type outbursts suggest that the $\alpha$ values in the inner disk region directly involved in the outburst can be much larger than that in the outer disk (e.g. Zhu et al., 2020). During the MRI-type outbursts, the inner disk gets significantly hot and exceeds ionization temperature for alkali metals of $\approx 1000$ K (Umebayashi & Nakano, 1981, 1988). The local thermal ionization then increases significantly and dominates the total ionization fraction. Thus, we set $\alpha_{\rm MRI}$ to the maximum value of 0.1, when the local thermal ionization fraction exceeds $10^{-10}$. The larger value of the $\alpha$ parameter results in a better fit to the duration of FUor-type outbursts that is inferred from observations (Zhu et al., 2009; Vorobyov et al., 2020b). The ionization threshold value is larger than that required for MRI activation in the midplane, however, it is justified, since the upper regions of the disk are significantly more ionized from external irradiation (Bai, 2011; Desch & Turner, 2015). The method of transitioning $\alpha_{\rm MRI}$ as a function of ionization fraction is a better approximation of disk physics, as compared to the previous criterion in Vorobyov et al. (2020b), where this switch was made for the entire innermost 5 au region. Inside the dead zone, a small value of residual viscosity corresponding to $\alpha_{{\rm dz}}=10^{-5}$ is considered, which may arise from purely hydrodynamic turbulence driven by the Maxwell stress in the active layer (Fleming & Stone, 2003; Okuzumi & Hirose, 2011).
@@
where $e$ is the charge of an electron, $m_{\rm e}$ is the mass of an electron and $\langle\sigma v\rangle_{\rm{en}}$ is the slowing-down coefficient (see Vorobyov et al., 2020b, for more details). We note that $\langle\sigma v\rangle_{\rm{en}}$ is now updated to the empirically verified $10^{-9}\,\mbox{cm}^{3}\,\mbox{s}^{-1}$ (Draine, 2011; Das & Basu, 2021), instead of the previously used value of $10^{-7}\,\mbox{cm}^{3}\,\mbox{s}^{-1}$ that was based on a classical estimate (Nakano, 1984). The dead zone is present only when $\Sigma_{\rm g}>2\Sigma_{\rm crit}$, in which case $\Sigma_{\rm MRI}$ is set to $2\Sigma_{\rm crit}$. If $\Sigma_{\rm g}\leq 2\Sigma_{\rm crit}$, the disk is considered to be fully MRI active and there is no dead zone. The MRI-active surface layer, as calculated with Eq. 12, depends sensitively on the ionization balance, thermal balance and the magnetic field evolution in the disk. However, it results in a low-viscosity dead zone in the innermost $\sim 10$ au of the disk (Eq. 10), as anticipated by simulations (Gammie, 1996; Dzyurkevich et al., 2010; Vorobyov et al., 2020b). The evolution of $B_{\rm z}$ is computed by solving the advection equation and the planar components of magnetic field, $B_{\rm p}$, is evaluated as the solution of a magnetic analogue of the Poisson integral (Vorobyov et al., 2020b). Note that the assumption of flux-freezing limit leads to significantly overestimated $B_{\rm z}$, which has consequences for the disk wind model.
@@ 2.2 Magnetic disk wind model
The magnetic disk wind model that we adopt is primarily based on the insights gained from the local MHD shearing box simulations of Bai (2013), which we modify for long-term disk evolution. The results of a series of shearing box simulations conducted at different radii can be stitched together to obtain a global picture of wind properties throughout the disk. Bai (2013) derived the following fitting formulae for the radial dependence of wind mass-loss rate (more accurately, rate of gas mass loss per unit area of the disk) and the wind-driven surface stress (i.e., the $z\phi$ component of Maxwell’s stress tensor) for a protoplanetary disk (his Eq. 9 and 10)
@@
Before Eqs. (13) and (14) could be used in a magnetohydrodynamic model, however, they need several modifications and here we will elaborate on these changes. Firstly, the radial term in these formulae is essentially an MMSN radius, in the sense that it corresponds to the stellar mass of 1 $M_{\odot}$ ($r\equiv r_{\rm M_{\odot}}$). A shearing box is assigned a radial location in the disk based on its average angular velocity, $\Omega$. Thus, to modify Eqs. (13) and (14) for an arbitrary stellar mass, the $(r_{\rm M_{\odot}}/{\rm au})$ term needs to be modified so as to have the same angular velocity
@@
${r_{\rm M_{\odot}}}=r\bigg{(}\frac{M_{\star}}{M_{\odot}}\bigg{)}^{-1/3},$ (16)
@@
Calculation of mass and angular momentum loss due to winds as prescribed by Eqs. (13) and (14) requires the disk plasma $\beta_{0}$, i.e., the midplane ratio of gas to magnetic pressure. Since FEOSAD evolves the magnetic field, $\beta_{0}$ can be obtained throughout the disk (Section 2.1). We note that $\beta_{0}$ is defined differently in Bai (2013) (Heaviside-Lorentz units, $\beta_{\rm wind}={2P_{0}}/{B_{z}^{2}}$) and in the hydrodynamic model of Vorobyov et al. (2020b) (Gaussian units, $\beta_{\rm MHD}={8\pi P_{0}}/{B_{z}^{2}}$), where $P_{0}$ and $B_{z}$ are midplane pressure and vertical magnetic field, respectively. In the context of calculating the wind quantities, $\beta_{0}=\beta_{\rm wind}$ is used. When the mass loss rate due to the winds over the entire disk is calculated by post-processing FEOSAD simulations (i.e., Eq. 13 applied to surface density of a gravitoviscous simulation), we found that this integrated mass loss rate was consistently about two orders of magnitude larger than the accretion rate onto the star. Such a large rate of mass loss is unrealistic and not sustainable, as the entire disk will be lost within $10^{4}$ yr.
There are two reasons for a substantially large wind mass loss rate while using Bai (2013)’s equations as they stand. Firstly, we found that the local $\beta_{0}$ in FEOSAD simulations is generally much lower than the values widely accepted in the literature. Measuring the magnetic field strength in protoplanetary disks remains challenging, although it is expected to be strongly subthermal, with a typical T Tauri disk exhibiting $\beta_{0}\lesssim 10^{4}$ (Vlemmings et al., 2019; Lesur, 2021b). Similarly large values of $\beta_{0}\sim 10^{2}-10^{8}$ are often considered in shearing box simulations (e.g. Bai, 2013; Simon et al., 2018), which are in alignment with global disk simulations that include non-ideal MHD effects (Masson et al., 2016; Xu & Kunz, 2021; Mauxion et al., 2024). However, in FEOSAD simulations, $\beta_{0}$ is consistently lower, indicating the presence of a strong magnetic field. This discrepancy occurs due to the ideal MHD limit for evolving the magnetic field, where the accreting gas tends to drag the field with it, increasing $B_{\rm z}$ and thus, decreasing local values of $\beta_{0}$. Secondly, the shearing box-simulations may severely overestimate the mass loss rate, as the mass loss rate consistently decreases with an increase in the box height in the direction normal to the disk midplane. This term is highly uncertain and Bai (2013) proposed an additional correction proportional to $H_{g}/r$, in order to account for this dependence on the finite size of the shearing box. In our model, we attempt to remedy the large magnetic fields in the MHD simulations via a constant correction factor,
@@
that is multiplied with $\beta_{0}$ while calculating the winds. This factor reflects the limitations of current simulations and may be dropped in the future when the non-ideal MHD equations are solved. Parameter $C_{\beta}$ is constrained such that $C_{\beta}\times\beta_{0}\sim 10^{4}$ over majority of the disk, which is close to the expected values of $\beta_{0}$ in protoplanetary disks. However, note that there is relatively large amount of uncertainty in this estimate from both observational and theoretical side. In order to account for the finite size of the shearing box, we include a term proportional to the local disk aspect ratio in the equation for wind mass loss $\dot{M}_{w}$
@@
where ${C_{H}}$ is a correction factor, with the default value of 0.5. When both ${\rm C_{\beta}}$ and ${\rm C_{Box}}$ terms are considered, the integrated disk wind mass loss rate is about the same order of magnitude as the accretion rate onto the star. The correction ${\rm C_{Box}}$ only applies to the mass loss rate and not the surface stress, as the latter is well constrained in the shearing box simulations. This parameter is constrained by the total mass loss rate from the winds, such that the latter is of the order of the accretion rate onto the central protostar (Pascucci et al., 2023). Concerning the total mass loss from the system, we additionally assume formation of a high velocity jet close to the star in all models, including those with only gravitoviscous evolution. The jet is formed concurrently with the star and it lies inside the central sink cell, and 10% of the accreting mass is assumed to be ejected via the jet (Hartigan et al., 1994; Bally, 2016).
@@
In the shearing box simulations that result in the relations specified in Eqs. (13) and (14), the FUV penetration depth is considered to be a constant ($\Sigma_{\rm FUV}=0.03$ ${\rm g\,cm^{-2}}$, in accordance with Perez-Becker & Chiang (2011)) and effects of changing $\Sigma_{\rm FUV}$ are not parameterized. To the first approximation, when $\Sigma_{\rm FUV}$ is increased by an order of magnitude, the wind mass loss increases by about the same factor, while the wind stress also increases, but only by a factor of few (Bai & Stone, 2013). This guides the calibration of the exponents in the following correction terms that are introduced, in order to account for varying FUV luminosity,

### v3

@@ 2.1 Disk evolution with FEOSAD
In this section we summarize the underlying MHD model that is used for the evolution of the PPD in the thin-disk limit. The magnetic disk winds described in Sect. 2.2 are added to this base model in a modular fashion. The hydrodynamic simulations are conducted using the code Formation and Evolution Of a Star And its circumstellar Disk (FEOSAD, Vorobyov et al., 2018a, 2020b). The numerical implementation of hydrodynamic algorithms is based on the operator-split procedure similar in methodology to the ZEUS code (Stone & Norman, 1992). We take into account the self-gravity of the disk, co-evolution of the coupled two-part dust component, calculation of magnetic fields in the flux-freezing approximation, and a dead zone modeled with an adaptive-$\alpha$ formulation, where the effective $\alpha$ is derived from ionization-recombination balance equations. Here we describe the minimum set of equations, while a more detailed description of the model can be obtained in Vorobyov et al. (2020b) and references therein.
The equations of mass continuity, conservation of momentum, and energy transport for the gas component in the thin disk limit can be described as follows:
@@
The energy transport equation (Eq. (3)) pertains to the evolution of the internal energy per surface area, $e$. The disk midplane temperature ($T_{\rm mp}$), which is assumed equal for the dust and gas components, is obtained by solving this energy balance equation. The heating ($\Gamma$) and cooling ($\Lambda$) rates are based on the analytical solution of the radiation transfer equations in the vertical direction (Dong et al., 2016). The heating function takes into account the irradiation of the disk surface from the stellar as well as background irradiation and radiation cooling by dust grains is also considered (Vorobyov et al., 2018a). The irradiation from the central star is calculated using incidence angle of radiation arriving at the disk surface, which in turn is obtained using the local scale height and a fixed flaring angle (Vorobyov & Basu, 2010b; Dong et al., 2016). These calculations neglect the self-shielding of the disk, which may arise in the innermost regions due to formation of concentric rings. However, this effect is considered to be insignificant, as viscous heating dominates the energy balance in the midplane of this inner disk region. The resulting model has a flared structure, wherein the disk vertical scale height increases with radial distance. Both the disk and the infalling envelope receive a fraction of the irradiation energy from the central protostar, both in terms of accretion and photospheric luminosity (Vorobyov & Basu, 2009). The optical depths in the calculations are proportional to the total dust surface density; however, they do not take into account the dust growth. The Planck and Rosseland mean opacities needed for these calculations are taken from Semenov et al. (2003). The term involving viscous stress tensor accounts for the viscous heating, which typically dominates the inner disk and becomes significant for an extended region during outbursts (Vorobyov & Basu, 2010b; Kadam et al., 2020). The last term in Eq. (3) corresponds to the internal energy of matter carried away by the magnetic winds.
@@
where $\Sigma_{\rm d,sm}$ and $\Sigma_{\rm d,gr}$ are the surface densities of small and grown dust, respectively. The term $u$ denotes the planar components of the grown dust velocity and $\dot{\Sigma}_{\rm w,d}$ accounts for the small dust blown away with the disk winds. Since the winds originate in the upper layers of the disk surface, we assume that they do not carry grown dust. $S$ is the rate of conversion from small to grown dust per unit surface area, which is a function of the maximum size of the dust $(a_{\rm max})$. The dust is assumed to mix vertically with the gas, which is a reasonable approximation for a young disk evolving under gravitational and viscous torques. $S(a_{\rm max})$ is derived from the assumption that the size distributions of both the dust populations can be described by a power law with an exponent of $-3.5$. The discontinuity in the dust size distribution at $a_{*}$ is assumed to get smoothed out, which implies the dominant role of dust growth as compared to the dust flow (Molyarova et al., 2021). The evolution of the maximum size of the grown dust can be expressed as an advection equation modified for the presence of a source term:
@@
where $v_{\rm frag}$ is the fragmentation velocity, $c_{\rm s}$ is the sound speed and $\alpha_{\rm turb}$ corresponds to the turbulence parameter (Birnstiel et al., 2012), wherein the possible contribution from gravitoturbulence is not considered (see, e.g., Vorobyov et al., 2023). We set the fragmentation velocity to a canonical value of 3 ${\rm m\,s^{-1}}$, although there is significant ambiguity due to sticking properties of different grain compositions, for example, bare silicate versus icy grains (Blum & Wurm, 2008; Wada et al., 2009; Steinpilz et al., 2019). When $a_{\rm max}$ exceeds $a_{\rm frag}$, the growth rate $\cal{D}$ is set to zero and $a_{\rm max}$ is set equal to $a_{\rm frag}$, which guarantees that the particle size is capped at $a_{\rm frag}$. We note that in our approach, the dust growth is limited to keep the size of dust particles within the Epstein regime so that the back-reaction term is expressed as
@@
The viscosity in a protoplanetary disk is thought to originate primarily due to the turbulence generated by MRI (Balbus & Hawley, 1991; Turner et al., 2014). As the cosmic rays externally penetrate the disk surface, the inner disk ($\lesssim 10$ au) accretes through the MRI-active surface layers, while a magnetically dead zone is formed at the disk midplane (Gammie, 1996; Simon et al., 2013). Although there may be quantitative uncertainty, this picture of layered accretion is considered generally valid, even in the presence of disk winds (Armitage & Kley, 2019). We modeled the magnetically dead zone using an effective and adaptive Shakura & Sunyaev (1973) $\alpha$ parameter (Bae et al., 2014; Kadam et al., 2019):
@@
where $\Sigma_{{\rm MRI}}$ is the gas surface density of the MRI-active layer and $\Sigma_{{\rm dz}}$ is that of the magnetically dead layer so that $\Sigma_{\rm g}=\Sigma_{{\rm MRI}}+\Sigma_{{\rm dz}}$. Here, $\alpha_{\rm MRI}$ and $\alpha_{{\rm dz}}$ correspond to the strength of the turbulence in the MRI-active layer and the dead zone, respectively. The $\Sigma_{\rm MRI}$ is obtained from the ionization fraction ($x$), which in turn is determined by solving the ionization balance equation
@@
where $\xi$ is the ionization rate, $\alpha_{\rm{r}}$ is radiative recombination rate, $n_{\rm n}$ is the number density of neutrals, and $\alpha_{\rm{d}}$ is the total rate of recombination on to the dust grains (Dudorov & Khaibrakhmanov, 2014; Dudorov & Sazonov, 1987; Balduin et al., 2023). We consider the ionization by cosmic rays and radionuclides; however, effects of stellar far-ultraviolet and X-ray radiation are neglected since their contribution is typically significantly smaller in the inner disk and limited to the upper disk atmosphere, which is not resolved in our thin-disk model (Igea & Glassgold, 1999; Bergin et al., 2007; Perez-Becker & Chiang, 2011).
In the simulations, $\alpha_{\rm MRI}$ is set to the canonical value of $0.01$, while lower values of $10^{-3}$ and $10^{-4}$ are also explored, the latter being consistent with recent observations of low viscosity in PPDs (Flaherty et al., 2015; Dullemond et al., 2018; Muro-Arena et al., 2018; Flaherty et al., 2020; Franceschi et al., 2023). 3D MHD simulations of FU Orionis type outbursts suggest that the $\alpha$ values in the inner disk region directly involved in the outburst can be much larger than that in the outer disk (e.g., Zhu et al., 2020). During the MRI-type outbursts, the inner disk gets significantly hot and exceeds ionization temperature for alkali metals of $\approx 1000$ K (Umebayashi & Nakano, 1981, 1988). The local thermal ionization then increases significantly and dominates the total ionization fraction. Thus, we set $\alpha_{\rm MRI}$ to the maximum value of 0.1, when the local thermal ionization fraction exceeds $10^{-10}$. The larger value of the $\alpha$ parameter results in a better fit to the duration of FUor-type outbursts that is inferred from observations (Zhu et al., 2009; Vorobyov et al., 2020b). The ionization threshold value is larger than that required for MRI activation in the midplane; however, it is justified since the upper regions of the disk are significantly more ionized from external irradiation (Bai, 2011; Desch & Turner, 2015). The method of transitioning $\alpha_{\rm MRI}$ as a function of ionization fraction is a better approximation of disk physics, as compared to the previous criterion in Vorobyov et al. (2020b), where this switch was made for the entire innermost five au region. Inside the dead zone, a small value of residual viscosity corresponding to $\alpha_{{\rm dz}}=10^{-5}$ is considered, which may arise from purely hydrodynamic turbulence driven by the Maxwell stress in the active layer (Fleming & Stone, 2003; Okuzumi & Hirose, 2011).
@@
where $e$ is the charge of an electron, $m_{\rm e}$ is the mass of an electron and $\langle\sigma v\rangle_{\rm{en}}$ is the slowing-down coefficient (see Vorobyov et al., 2020b, for more details). We note that $\langle\sigma v\rangle_{\rm{en}}$ is now updated to the empirically verified $10^{-9}\,\mbox{cm}^{3}\,\mbox{s}^{-1}$ (Draine, 2011; Das & Basu, 2021), instead of the previously used value of $10^{-7}\,\mbox{cm}^{3}\,\mbox{s}^{-1}$ that was based on a classical estimate (Nakano, 1984). The dead zone is present only when $\Sigma_{\rm g}>2\Sigma_{\rm crit}$, in which case $\Sigma_{\rm MRI}$ is set to $2\Sigma_{\rm crit}$. If $\Sigma_{\rm g}\leq 2\Sigma_{\rm crit}$, the disk is considered to be fully MRI active and there is no dead zone. The MRI-active surface layer, as calculated with Eq. 12, depends sensitively on the ionization balance, thermal balance and the magnetic field evolution in the disk. However, it results in a low-viscosity dead zone in the approximately innermost ten au of the disk (Eq. 10), as anticipated by simulations (Gammie, 1996; Dzyurkevich et al., 2010; Vorobyov et al., 2020b). The evolution of $B_{\rm z}$ is computed by solving the advection equation and the planar components of magnetic field, $B_{\rm p}$, is evaluated as the solution of a magnetic analogue of the Poisson integral (Vorobyov et al., 2020b). We note that the assumption of flux-freezing limit leads to significantly overestimated $B_{\rm z}$, which has consequences for the disk wind model.
@@ 2.2 Magnetic disk wind model
The magnetic disk wind model that we adopt is primarily based on the insights gained from the local MHD shearing box simulations of Bai (2013), which we modify for long-term disk evolution. The results of a series of shearing box simulations conducted at different radii can be stitched together to obtain a global picture of wind properties throughout the disk. Bai (2013) derived the following fitting formulae for the radial dependence of wind mass loss rate (more accurately, rate of gas mass loss per unit area of the disk) and the wind-driven surface stress (i.e., the $z\phi$ component of Maxwell’s stress tensor) for a protoplanetary disk (his Eqs. 9 and 10):
@@
Before Eqs. (13) and (14) could be used in a magnetohydrodynamic model, however, they need several modifications, and here we elaborate on these changes. Firstly, the radial term in these formulae is essentially an MMSN radius, in the sense that it corresponds to the stellar mass of one $M_{\odot}$ ($r\equiv r_{\rm M_{\odot}}$). A shearing box is assigned a radial location in the disk based on its average angular velocity, $\Omega$. Thus, to modify Eqs. (13) and (14) for an arbitrary stellar mass, the $(r_{\rm M_{\odot}}/{\rm au})$ term needs to be modified so as to have the same angular velocity:
@@
${r_{\rm M_{\odot}}}=r\bigg{(}\frac{M_{\star}}{M_{\odot}}\bigg{)}^{-1/3}$ (16)
@@
Calculation of mass and angular momentum loss due to winds as prescribed by Eqs. (13) and (14) requires the disk plasma $\beta_{0}$, that is, the midplane ratio of gas to magnetic pressure. Since FEOSAD evolves the magnetic field, $\beta_{0}$ can be obtained throughout the disk (Sect. 2.1). We note that $\beta_{0}$ is defined differently in Bai (2013) (Heaviside-Lorentz units, $\beta_{\rm wind}={2P_{0}}/{B_{z}^{2}}$) and in the hydrodynamic model of Vorobyov et al. (2020b) (Gaussian units, $\beta_{\rm MHD}={8\pi P_{0}}/{B_{z}^{2}}$), where $P_{0}$ and $B_{z}$ are midplane pressure and vertical magnetic field, respectively. In the context of calculating the wind quantities, $\beta_{0}=\beta_{\rm wind}$ is used. When the mass loss rate due to the winds over the entire disk is calculated by post-processing FEOSAD simulations (i.e., Eq. 13 applied to surface density of a gravitoviscous simulation), we found that this integrated mass loss rate was consistently about two orders of magnitude larger than the accretion rate onto the star. Such a high rate of mass loss is unrealistic and not sustainable, as the entire disk will be lost within $10^{4}$ yr.
There are two reasons for a substantially large wind mass loss rate while using Bai (2013)’s equations as they stand. Firstly, we found that the local $\beta_{0}$ in FEOSAD simulations is generally much lower than the values widely accepted in the literature. Measuring the magnetic field strength in protoplanetary disks remains challenging, although it is expected to be strongly subthermal, with a typical T Tauri disk exhibiting $\beta_{0}\lesssim 10^{4}$ (Vlemmings et al., 2019; Lesur, 2021b). Similarly large values of $\beta_{0}\sim 10^{2}-10^{8}$ are often considered in shearing box simulations (e.g., Bai, 2013; Simon et al., 2018), which are in alignment with global disk simulations that include non-ideal MHD effects (Masson et al., 2016; Xu & Kunz, 2021; Mauxion et al., 2024). However, in FEOSAD simulations, $\beta_{0}$ is consistently lower, indicating the presence of a strong magnetic field. This discrepancy occurs due to the ideal MHD limit for evolving the magnetic field, where the accreting gas tends to drag the field with it, increasing $B_{\rm z}$ and thus decreasing local values of $\beta_{0}$. Secondly, the shearing box-simulations may severely overestimate the mass loss rate, as the mass loss rate consistently decreases with an increase in the box height in the direction normal to the disk midplane. This term is highly uncertain and Bai (2013) proposed an additional correction proportional to $H_{g}/r$, in order to account for this dependence on the finite size of the shearing box. In our model, we attempt to remedy the large magnetic fields in the MHD simulations via a constant correction factor,
@@
that is multiplied with $\beta_{0}$ while calculating the winds. This factor reflects the limitations of current simulations and may be dropped in the future when the non-ideal MHD equations are solved. Parameter $C_{\beta}$ is constrained such that $C_{\beta}\times\beta_{0}\sim 10^{4}$ over majority of the disk, which is close to the expected values of $\beta_{0}$ in protoplanetary disks. However, we note that there is relatively large amount of uncertainty in this estimate from both observational and theoretical side. In order to account for the finite size of the shearing box, we include a term proportional to the local disk aspect ratio in the equation for wind mass loss $\dot{M}_{w}$:
@@
where ${C_{H}}$ is a correction factor with the default value of 0.5. When both the ${\rm C_{\beta}}$ and ${\rm C_{Box}}$ terms are considered, the integrated disk wind mass loss rate is about the same order of magnitude as the accretion rate onto the star. The correction ${\rm C_{Box}}$ only applies to the mass loss rate and not the surface stress, as the latter is well constrained in the shearing box simulations. This parameter is constrained by the total mass loss rate from the winds such that the latter is of the order of the accretion rate onto the central protostar (Pascucci et al., 2023). Concerning the total mass loss from the system, we additionally assume formation of a high velocity jet close to the star in all models, including those with only gravitoviscous evolution. The jet is formed concurrently with the star and it lies inside the central sink cell, and 10% of the accreting mass is assumed to be ejected via the jet (Hartigan et al., 1994; Bally, 2016).
@@
In the shearing box simulations that result in the relations specified in Eqs. (13) and (14), the FUV penetration depth is considered to be a constant ($\Sigma_{\rm FUV}=0.03$ ${\rm g\,cm^{-2}}$, in accordance with Perez-Becker & Chiang (2011)) and effects of changing $\Sigma_{\rm FUV}$ are not parameterized. To the first approximation, when $\Sigma_{\rm FUV}$ is increased by an order of magnitude, the wind mass loss increases by about the same factor, while the wind stress also increases, but only by a factor of a few (Bai & Stone, 2013). This guides the calibration of the exponents in the following correction terms that are introduced in order to account for varying FUV luminosityin the wind mass loss rate and surface stress equations, respectively: